Average Rate of Return Method, Net Present Value Method, and Analysis for a service company

The capital investment committee of Arches Landscaping Company is considering two capital investments. The estimated income from operations and net cash flows from each investment are as follows:

Front-End Loader Greenhouse
Year Income from
Operations
Net Cash
Flow
Income from
Operations
Net Cash
Flow
1 $45,600 $141,000 $96,000 $226,000
2 45,600 141,000 73,000 190,000
3 45,600 141,000 36,000 134,000
4 45,600 141,000 16,000 92,000
5 45,600 141,000 7,000 63,000
Total $228,000 $705,000 $228,000 $705,000

Each project requires an investment of $480,000. Straight-line depreciation will be used, and no residual value is expected. The committee has selected a rate of 10% for purposes of the net present value analysis.

Present Value of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 0.890 0.826 0.797 0.756 0.694
3 0.840 0.751 0.712 0.658 0.579
4 0.792 0.683 0.636 0.572 0.482
5 0.747 0.621 0.567 0.497 0.402
6 0.705 0.564 0.507 0.432 0.335
7 0.665 0.513 0.452 0.376 0.279
8 0.627 0.467 0.404 0.327 0.233
9 0.592 0.424 0.361 0.284 0.194
10 0.558 0.386 0.322 0.247 0.162

Required:

1a. Compute the average rate of return for each investment. If required, round your answer to one decimal place.

Average Rate of Return
Front-End Loader %
Greenhouse %

1b. Compute the net present value for each investment. Use the present value of $1 table above. If required, use the minus sign to indicate a negative net present value.

Front-End Loader Greenhouse
Present value of net cash flow $ $
Amount to be invested $ $
Net present value $ $
Answer #1
Average net income 45600 =228000/5
Average Investment 240000 =480000/2
1a
Average Rate of Return
Front-End Loader 19% =45600/240000
Greenhouse 19% =45600/240000
1b
Front-End Loader Greenhouse
Present value of net cash flow 534390 564967
Amount to be invested 480000 480000
Net present value 54390 84967
Workings:
Front-End Loader Greenhouse
Cash flows PV factor 10% Present value Cash flows PV factor 10% Present value
1 141000 0.909 128169 226000 0.909 205434
2 141000 0.826 116466 190000 0.826 156940
3 141000 0.751 105891 134000 0.751 100634
4 141000 0.683 96303 92000 0.683 62836
5 141000 0.621 87561 63000 0.621 39123
Total 534390 Total 564967